# Beyond Diagonal RIS for 6G Non-Terrestrial Networks: Potentials and Challenges

### v1

@@ III-A1 Reflective BD-RIS
In this mode, the BD-RIS reflects signals that come from one side back towards that side, resulting in coverage of 180-degree space. As can be seen in the satellite communication scenario in Fig. 2(i), all of the reflecting elements of the BD-RIS are oriented in the same direction in order to enable the reflective mode during transmission. When considering the BD-RIS with reflecting mode, the phase shift matrix $\boldsymbol{\Phi}$, which is subject to a unitary constraint, stands out as the defining characteristic. The matrix $\boldsymbol{\Phi}$ is unitary if it satisfies the following condition: $\boldsymbol{\Phi}^{*}\boldsymbol{\Phi}=\boldsymbol{\Phi}\boldsymbol{\Phi}^{*}={\bf I}$, where $\boldsymbol{\Phi}^{*}$ is the conjugate transpose of $\boldsymbol{\Phi}$ and ${\bf I}$ denotes an identity matrix. This property ensures that the columns (and rows) of the matrix are orthogonal. In the context of BD-RIS for 6G wireless networks, unitary constraints are applied to the design of these surfaces to ensure efficient signal manipulation without loss of power. For example, the BD-RIS may be designed to apply unitary transformations to incoming signals, maintaining signal integrity and improving communication performance by optimizing reflection and transmission paths.
III-A2 Hybrid BD-RIS
@@
This mode allows signals that are impinging on one side of the BD-RIS to be partially reflected toward the same side and partially transmitted toward the opposing side, which results in a coverage of 360-degree space. The BD-RIS that operates in hybrid mode is also referred to as the intelligent omni-surface or the simultaneous transmitting and reflecting BD-RIS. In order to support the hybrid mode, two elements with a unidirectional radiation pattern are placed back to back. These elements are then connected to a two-port fully connected reconfigurable impedance network, as shown in HAPS communication scenario in Fig. 2(ii). This ensures that each reflecting element and each transmitting element cover 180-degree space, respectively, in order to achieve 360-degree coverage. In terms of mathematics, the BD-RIS with $K$ number of phase response elements and operating in hybrid mode is characterized by two matrices, namely transmitting and reflecting, i.e., $\boldsymbol{\Phi}_{t}\in\mathbb{C}^{K/2\times K/2}$ and $\boldsymbol{\Phi}_{r}\in\mathbb{C}^{K/2\times K/2}$, where $\boldsymbol{\Phi}_{t}$ is a matrix of transmitting phase response elements and $\boldsymbol{\Phi}_{r}$ denotes the matrix of reflecting phase response elements. These matrices should satisfy $\boldsymbol{\Phi}^{H}_{t}\boldsymbol{\Phi}_{t}+\boldsymbol{\Phi}^{H}_{r}\boldsymbol{\Phi}_{r}={\bf I}_{K/2}$.
III-A3 Multi-Sector BD-RIS
@@
This mode is an extension or a more comprehensive version of the hybrid mode. In multi-sector BD-RIS, the entire area is partitioned into $S$ sectors, where $S$ must be greater than or equal to 2. Signals that come into contact with one sector of the BD-RIS can be partially redirected back to the same sector and partially dispersed toward the remaining $S-1$ sectors. In order to facilitate the multi-sector mode of BD-RIS, each sector consists of $S$ phase response elements positioned at the edges of an $S$-sided polygon. These elements have a unidirectional radiation pattern that covers the $1/2$ area of the sector, preventing any overlap between adjacent sectors. The $S$ elements are connected to a $S$-port fully connected reconfigurable impedance network, as illustrated in UAV communication scenario in 2(iii). Therefore, the multi-sector mode of BD-RIS can cover 360-degree space, similar to the hybrid BD-RIS. However, multi-sector BD-RIS offers greater performance improvements compared to the hybrid mode due to the utilization of high-gain elements with narrower beamwidth, which covers 180-degree space. The BD-RIS with multi-sector mode can be mathematically described by a set of $S$ matrices, $\boldsymbol{\Phi}_{s}\in\mathbb{C}^{K/S\times K/S}$, $s\in S$. These matrices should meet the condition of $\sum_{s=1}^{S}\boldsymbol{\Phi}_{s}^{H}\boldsymbol{\Phi}_{s}={\bf I}_{K/S}$.
Figure 2: Different modes of BD-RIS.
III-B Advantages of BD-RIS in NTNs
BD-RIS provides substantial benefits compared to the traditional diagonal RIS system due to the connection between elements and the full scattering matrix. BD-RIS offers advanced beamforming capabilities, superior signal manipulation, enhanced flexibility and adaptability, increased spectrum efficiency, improved interference management, and greater resilience to environmental changes. These advances greatly enhance the capabilities of BD-IRS, making it a highly potent and adaptable technology for improving future wireless communication networks, particularly those designed for 6G and beyond. Key advantages of BD-RIS over diagonal RIS can be further discussed in the following.
III-B1 Advance Beamforming
@@
The functionality of conventional RIS is restricted to controlling the phase of the incident signal. As a result, its capacity to manipulate passive beamforming is limited, leading to a degradation in performance. BD-RIS provides inter-element connections, however, this comes at the cost of increased circuit complexity; its mathematical description is not restricted to diagonal phase response matrices. This makes for more precise and adaptive beamforming, which boosts signal quality and strength in many directions.
III-B2 Improved Signal Manipulation
In contrast to the conventional diagonal RIS, which can only manipulate the diagonal elements of the phase response matrix, BD-RIS comparatively offers greater flexibility by allowing manipulation of both diagonal and non-diagonal elements of the phase response matrix. This enhanced capability significantly improves the performance of various wireless communication scenarios in 6G networks.
III-B3 Enhanced Flexibility and Adaptability
Conventional RIS is efficient in situations where signal paths are more consistent and can be predicted. BD-RIS, in contrast, offers enhanced adaptability to accommodate fluctuating signal conditions, including movements in user positions, mobile blockages, and evolving network requirements. This flexibility guarantees constant performance even under extremely unpredictable settings.
III-B4 Better Interference Management
Diagonal RIS can reduce interference by adjusting the phase of incident signals. BD-IRS offers exceptional interference management capabilities through the utilization of modern signal processing techniques. This leads to reduced levels of interference, hence improving the overall performance of the network and the experience of the user.
III-B5 360-Degree Coverage
Unlike diagonal RIS, which has a limited coverage of 180-degrees, the BD-RIS operates in hybrid and multi-sector modes, providing 360-degree coverage. Furthermore, the multi-sector mode can enhance channel gain by utilizing smaller beamwidth and higher element gain of each BD-RIS element. This effectively extends the communication range, ensuring full-space coverage.
@@ IV Applications of BD-RIS in NTNs
Well-established research on conventional diagonal RIS, clearly demonstrates its use cases and applications in diverse fields, including, but not limited to, wireless power transfer, coverage improvements, and wireless sensing. In the following sections, we showcase, for the first time, some of the potential applications of BD-RIS in an NTN-enabled communication scenario.
@@ IV-A Seamless and Ultra Precision Indoor Positioning
It is well-known that wireless signals based on NTN are unable to penetrate indoor environments effectively, as they are weakened and scattered by roofs, walls, and other structures. Moreover, the availability of line-of-sight (LoS) signals greatly impacts positioning accuracy. These limitations necessitate a separate infrastructure for indoor positioning applications, thereby increasing overall costs and energy requirements. Since BD-RIS can flexibly reflect signals in multiple controlled directions, it can play a pivotal role in providing an effective conduit for NTN signals to overcome penetration losses caused by building structures. Furthermore, compared to conventional diagonal RIS, BD-RIS offers enhanced beamforming capabilities, thus improving a network’s ability to achieve ultra-precise localization of devices. In conclusion, these additional features of BD-RIS can be crucial in developing a new class of indoor positioning systems that do not require any additional infrastructure, costs, or energy resources.
@@ IV-B Improved IoT Connectivity
The Internet of Things (IoT) has been a familiar concept for quite some time, but its popularity surged with the integration of machine-type communication (mMTC) as a feature of 5G. IoT encompasses a diverse range of applications, spanning from smart home automation to monitoring deforestation activities. Particularly in areas with limited or no cellular coverage, connecting these IoT devices to the internet poses a significant challenge. To address such obstacles, the 3GPP release 17 incorporates NTN into traditional Terrestrial Networks (TN), enabling NTNs to serve as a bridge between the internet and these IoT devices. Despite playing a critical role in overcoming these limitations, NTN introduces its own set of challenges, such as coverage (while considering smart home indoor environment applications in remote/rural areas) and power consumption. Similarly, the enhanced capabilities of BD-RIS, which include reflecting signals in multiple controlled directions and improved beamforming, have the potential to mitigate these barriers and establish a seamless connection between NTN and IoT devices.
@@ IV-C Extended Coverage for NTN-aided Terrestrial Users
One of the core applications of conventional diagonal RIS, for which it was originally developed, is to provide extended coverage, overcome blind spots, and ensure seamless connectivity for terrestrial users. In the context of NTN-aided communication, conventional diagonal RIS has shown great potential in improving the performance of cellular users with weak or no connectivity to the base station. For instance, as demonstrated in [14], conventional diagonal RIS-aided UAVs (a specific use case of NTN) have clearly illustrated the usefulness of conventional diagonal RIS in NTN systems. Although conventional diagonal RIS-aided NTN has shown significant potential in enhancing the performance of terrestrial users, the limitations associated with conventional diagonal RIS prevent us from fully benefiting from NTN. As discussed in the preceding sections, BD-RIS clearly surpasses conventional diagonal RIS by offering greater flexibility in reflecting signals in multiple controlled directions. This capability provides more extended coverage compared to conventional diagonal RIS and plays a significant role in increasing the performance of NTN.
Figure 3: BD-RIS enabled LEO communication.
V Case Study: BD-RIS Enabled NOMA LEO Satellite Communications
In this section, we provide a case study on BD-RIS enabled NTNs. We will first discuss the considered system model and proposed optimization. Then, we provide the numerical results and their discussion.
@@ V-A System Model and Proposed Optimization
Let us consider a downlink scenario of BD-RIS enabled satellite communication, as shown in Fig. 3, where an LEO communicates with ground users (GUs) using the power domain NOMA. In practice, the GUs operate with low directional antennas, so closing links between LEO and GUs is challenging when the network operates on a high frequency. Moreover, the direct links between LEO and GUs might not be possible, considering the mobile scenario. To compensate for path losses due to poor link budget and improve the received signal strength, we consider BD-RIS installed on the optimal location and assisting the signal delivery from LEO to GUs. Thus, GUs can receive signals from direct and BD-RIS enabled links. Assuming perfect acquisition of channel state information (CSI) by channel estimation techniques, this scenario provides an upper-bound solution.
This case study seeks to improve the achievable sum-rate of BD-RIS enabled NOMA LEO satellite communication by optimizing the transmission power of LEO and phase response of BD-RIS while ensuring the minimum data rate requirements of GUs. This is achieved by formulating a mathematical optimization problem for achievable sum-rate maximization which is subjected to several practical constraints such as minimum rate constraint, LEO power budget constraint, NOMA power allocation constraint, and BD-RIS phase response constraint, respectively. The formulated optimization is nonlinear and NP-hard in nature due to coupling decision variables and NOMA interference. In practice, solving such problems is challenging due to high computational complexity. To make the optimization tractable, we first adopt the block descent coordinate method to split the joint optimization into two problems: 1) transmission power at LEO for a given phase response at BD-RIS, and 2) phase response at BD-RIS, given the power allocation for LEO. Then, each problem is further transformed into linear form and simultaneously solved using a standard convex optimization such as CVX toolbox.
Figure 4: Performance Analysis

### v2

@@ III-B1 Reflective BD-RIS
In this mode, the feed antenna (transmitter) and receiver are located on the same side of BD-RIS, resulting in coverage of 180-degree space [^3] Transmissive BD-RIS, while similar to reflective BD-RIS, refracts signals instead of reflecting them, allowing the signals to pass through surfaces, as depicted in Fig. 3 (c). This mechanism is helpful for signal penetration through barriers like windows, walls, or partitions. Due to space limitations, we do not discuss transmissive BD-RIS here in detail.. The incident signal and the reflected signal occur on the same side of the surface, and self-interference results that impact the system performance [3]. As can be seen in Fig. 3(a), all of the PREs of the BD-RIS are oriented in the same direction in order to enable the reflective mode during transmission. When considering the BD-RIS with reflecting mode, the PRM $\boldsymbol{\Phi}$, which is subject to a unitary constraint, stands out as the defining characteristic. Since this BD-RIS mode can only reflect the signal, therefore $\boldsymbol{\Phi}_{r}$ is the unitary matrix which satisfies $\boldsymbol{\Phi}_{r}^{H}\boldsymbol{\Phi}_{r}={\bf I}_{K}$, where $\boldsymbol{\Phi}_{r}^{H}$ is the conjugate transpose of $\boldsymbol{\Phi}_{r}$ and ${\bf I}_{K}$ denotes an identity matrix. This property ensures that the columns (and rows) of the matrix are orthogonal. In the context of BD-RIS for 6G wireless networks, unitary constraints are applied to the design of these surfaces to ensure efficient signal manipulation without loss of power. Fig. 3(b) is an example of element wise reflective BD-RIS with group-connected architecture, where the number of PREs in each group is 4. In this example, the number of groups is $G=2$, where each group consists of $\hat{K}=K/G=4$ elements. The PRM can be defined as $\boldsymbol{\Phi}_{r}=\text{blkdiag}(\boldsymbol{\Phi}_{r,1},\boldsymbol{\Phi}_{r,2})$ where $\boldsymbol{\Phi}_{r,1}$ and $\boldsymbol{\Phi}_{r,2}$ satisfy $\boldsymbol{\Phi}_{r,1}^{H}\boldsymbol{\Phi}_{r,1}={\bf I}_{4}$ and $\boldsymbol{\Phi}_{r,2}^{H}\boldsymbol{\Phi}_{r,2}={\bf I}_{4}$. The reflective BD-RIS can be placed on walls, building facades, or ceilings to enhance signal coverage by bouncing the signal from the source to the destination.
III-B2 Hybrid BD-RIS
@@
This mode allows signals that are impinging on one side of the BD-RIS to be partially reflected toward the same side and partially transmitted toward the opposing side, which results in a coverage of 360-degree space. The BD-RIS that operates in single-connected hybrid mode is also referred to as the intelligent omni-surface or the simultaneous transmitting and reflecting BD-RIS. In order to support the hybrid mode, two elements with a unidirectional radiation pattern are placed back to back to form a cell. These elements are then connected to a two-port fully connected reconfigurable impedance network, as shown in Fig. 3(d). This ensures that each reflecting element and each transmitting element cover 180-degree space, respectively, in order to achieve 360-degree coverage. In terms of mathematics, a $K$-cell BD-RIS operating in hybrid mode is characterized by two matrices, namely reflecting and transmitting, i.e., $\boldsymbol{\Phi}_{r}$ and $\boldsymbol{\Phi}_{t}$. These matrices should satisfy $\boldsymbol{\Phi}^{H}_{r}\boldsymbol{\Phi}_{r}+\boldsymbol{\Phi}^{H}_{t}\boldsymbol{\Phi}_{t}={\bf I}_{K}$. Fig. 3(e) provides an example of cell-wise group-connected hybrid BD-RIS with 2 groups and 4 cells, where the group size is $\hat{K}=K/G=8/2=4$ and cell size is 2. In this case, $\boldsymbol{\Phi}_{r}$ and $\boldsymbol{\Phi}_{t}$ can be derived as $\boldsymbol{\Phi}_{r}=\text{blkdiag}(\boldsymbol{\Phi}_{r,1},\boldsymbol{\Phi}_{r,2})$ and $\boldsymbol{\Phi}_{t}=\text{blkdiag}(\boldsymbol{\Phi}_{t,1},\boldsymbol{\Phi}_{t,2})$. These matrices should satisfy as $\boldsymbol{\Phi}_{r,1}^{H}\boldsymbol{\Phi}_{r,1}+\boldsymbol{\Phi}_{t,1}^{H}\boldsymbol{\Phi}_{t,1}={\bf I}_{4}$ and $\boldsymbol{\Phi}_{r,2}^{H}\boldsymbol{\Phi}_{r,2}+\boldsymbol{\Phi}_{t,2}^{H}\boldsymbol{\Phi}_{t,2}={\bf I}_{4}$.
TABLE II: Characteristics of different BD-RIS architectures and modes.
Characteristics Single-connected architecture Fully-connected architecture Group-connected architecture
Number of groups $K$ 1 $G$
Group dimension 1 $K$ $\hat{K}$
Elements/group 1 $K^{2}$ $\hat{K}^{2}$
Number of Non-zero elements $K$ $K^{2}$ $G\hat{K}^{2}$
Reflective mode $|\phi_{r,k}|^{2}=1$ ${\bf\Phi}^{H}_{r}{\bf\Phi}_{r}={\bf I}_{K}$ ${\bf\Phi}^{H}_{r,g}{\bf\Phi}_{r,g}={\bf I}_{\hat{K}}$
Transmissive mode $|\phi_{t,k}|^{2}=1$ ${\bf\Phi}^{H}_{t}{\bf\Phi}_{t}={\bf I}_{K}$ ${\bf\Phi}^{H}_{t,g}{\bf\Phi}_{t,g}={\bf I}_{\hat{K}}$
Hybrid mode $|\phi_{r,k}|^{2}+|\phi_{t,k}|^{2}=1$ ${\bf\Phi}^{H}_{r}{\bf\Phi}_{r}+{\bf\Phi}^{H}_{t}{\bf\Phi}_{t}={\bf I}_{K}$ ${\bf\Phi}^{H}_{r,g}{\bf\Phi}_{r,g}+{\bf\Phi}^{H}_{t,g}{\bf\Phi}_{t,g}={\bf I}_{\hat{K}}$
Hardware complexity R: $K$, H: $(3/2)K$, M: $(S+1)K/2$ R,H,M: $(K+1)\frac{K}{2}$ R,H,M: $(\frac{K}{G}+1)\frac{K}{2}$
R-Reflective, H-Hybrid, M-Multi sector.
III-B3 Multi-Sector BD-RIS
@@
This mode is an extension or a more comprehensive version of the hybrid mode. In multi-sector BD-RIS, the entire area is partitioned into multiple sectors, where the number of sectors must be greater than or equal to 2, as shown in 3(f). Signals that come into contact with one sector of the BD-RIS can be partially redirected back to the same sector and partially dispersed toward the remaining sectors. Let us assume that $S$ is the number of sectors that constitute multi-sector BD-RIS; then, to facilitate the multi-sector mode of BD-RIS, each cell consists of $S$ PREs positioned at the edges of an $S$-sided polygon. These PREs have a unidirectional radiation pattern such that each PRE can cover the $1/S$ area of the sector, preventing any overlap between adjacent sectors. The $S$ PREs are connected to a $S$-port fully connected reconfigurable impedance network. Therefore, the multi-sector mode of BD-RIS can cover 360-degree space, similar to the hybrid BD-RIS. However, multi-sector BD-RIS offers greater performance improvements compared to the hybrid mode due to the utilization of high-gain PREs with narrower beamwidth, which can cover 180-degree space. The BD-RIS with multi-sector mode can be mathematically described by a set of $S$ matrices, $\boldsymbol{\Phi}_{s}\in\mathbb{C}^{K/S\times K/S}$, $s=\{1,2,\dots,S\}$, which fulfilling $\sum_{s=1}^{S}\boldsymbol{\Phi}_{s}^{H}\boldsymbol{\Phi}_{s}={\bf I}_{K/S}$. To further understand the concept of this mode, Fig. 3(g) provides cell-wise single-connected multi-sector BD-RIS with 3 sectors and 2 cells, where the number of BD-RIS elements in each cell is 3. The PRMs should satisfy as $\boldsymbol{\Phi}_{1}^{H}\boldsymbol{\Phi}_{1}+\boldsymbol{\Phi}_{2}^{H}\boldsymbol{\Phi}_{2}+\boldsymbol{\Phi}_{3}^{H}\boldsymbol{\Phi}_{3}={\bf I}_{2}$.
The detailed characteristics of different BD-RIS hardware architectures and operating modes are summarised in Table II.
III-C Advantages of BD-RIS in NTNs
BD-RIS provides substantial benefits compared to its conventional counterpart due to the connection between elements and the full scattering matrix. Some of the key advantages of BD-RIS over conventional RIS are discussed in the following:
III-C1 Advance BF and High Flexibility
@@
The functionality of conventional RIS is restricted to controlling the phase of the incident signal. As a result, its capacity to manipulate passive BF is limited, leading to a degradation in performance. BD-RIS provides inter-element connections; however, this comes at the cost of increased circuit complexity; its mathematical description is not restricted to diagonal PRMs. This makes for more precise and adaptive BF, which boosts signal quality and strength in many directions. Moreover, the flexibility of BD-RIS for manipulating diagonal and non-diagonal elements of the PRM significantly improves the system performance of various communication scenarios in 6G NTNs.
III-C2 Enhanced Adaptability to Challenging Environments
Due to its limited BF capability, CD-RIS is only effective in non-terrestrial scenarios with relatively stable and predictable links. In contrast, BD-RIS significantly enhances adaptability to more dynamic and challenging environments in NTNs. BD-RIS dynamically adjusts its configuration to accommodate changes in user positions. This ensures optimal signal reception and transmission even as users move, enhancing overall connectivity and reducing signal degradation. Moreover, in environments with moving obstacles or dynamic blockages, BD-RIS can effectively counteract signal interference caused by these obstructions. Its advanced reconfiguration capabilities allow it to maintain reliable communication links despite mobile barriers. Furthermore, the design of BD-RIS adapts to changing network demands, including varying traffic loads and shifting service requirements. This flexibility ensures consistent performance across different scenarios, accommodating both peak and off-peak traffic conditions.
III-C3 Better Interference Management
Unlike CD-RIS, which offers limited interference mitigation by adjusting the phase of incident signals, BD-RIS provides significantly enhanced interference management capabilities, particularly in NTNs. More specifically, BD-RIS employs advanced signal processing techniques and sophisticated configurations to manage interference more effectively than its conventional counterpart. In NTNs, where signal paths can be complicated because of satellites and other high-altitude platforms, BD-RIS works very well because it changes both the phase and amplitude of signals as they come in. This capability allows BD-RIS to adapt to varying interference conditions and mitigate interference from multiple sources, including both terrestrial and non-terrestrial transmissions.
III-C4 360-Degree Coverage
BD-RIS provides complete 360-degree coverage, unlike CD-RIS, which typically offers limited coverage of up to 180 degrees. BD-RIS operates in hybrid and multi-sector modes, allowing it to extend its coverage range effectively. The multi-sector mode leverages smaller beam widths and the higher element gain of each BD-RIS element to enhance channel gain. This approach not only broadens the communication range but also ensures full-space coverage, making BD-RIS ideal for environments requiring extensive and uniform signal distribution. The ability to deliver 360-degree coverage represents a significant advancement over CD-RIS configurations, providing more reliable and widespread communication in various NTN settings.
@@ IV Applications of BD-RIS in NTNs
Well-established research on CD-RIS clearly demonstrates its use cases and applications in diverse fields, including, but not limited to, wireless power transfer, coverage improvements, and wireless sensing. In the following sections, we showcase, for the first time, some of the potential applications of BD-RIS in an NTN-enabled communication scenario.
@@ IV-A Seamless and Ultra Precision Indoor Positioning
It is well-known that wireless signals based on NTN are unable to penetrate indoor environments effectively, as they are weakened and scattered by roofs, walls, and other structures. Moreover, the availability of line-of-sight (LoS) signals greatly impacts positioning accuracy. These limitations necessitate a separate infrastructure for indoor positioning applications, thereby increasing overall costs and energy requirements. Since BD-RIS can flexibly reflect signals in multiple controlled directions, it can play a pivotal role in providing an effective conduit for NTN signals to overcome penetration losses caused by building structures. Furthermore, compared to CD-RIS, BD-RIS offers enhanced BF capabilities, thus improving a network’s ability to achieve ultra-precise localization of devices. In conclusion, these additional features of BD-RIS can be crucial in developing a new class of indoor positioning systems that do not require any additional infrastructure, costs, or energy resources.
@@ IV-B Improved IoT Connectivity
The Internet of Things (IoT) has been a familiar concept for quite some time, but its popularity surged with the integration of machine-type communication (mMTC) as a feature of 5G. IoT encompasses a diverse range of applications, spanning from smart home automation to monitoring deforestation activities. Particularly in areas with limited or no cellular coverage, connecting these IoT devices to the internet poses a significant challenge. To address such obstacles, the 3GPP release 17 incorporates NTN into traditional Terrestrial Networks (TN), enabling NTNs to serve as a bridge between the internet and these IoT devices. Despite playing a critical role in overcoming these limitations, NTN introduces its own set of challenges, such as coverage (while considering smart home indoor environment applications in remote/rural areas) and power consumption. Similarly, the enhanced capabilities of BD-RIS, which include reflecting signals in multiple controlled directions and improved BF, have the potential to mitigate these barriers and establish a seamless connection between NTN and IoT devices.
@@ IV-C Extended Coverage for NTN-aided Terrestrial Users
One of the core applications of CD-RIS, for which it was originally developed, is to provide extended coverage, overcome blind spots, and ensure seamless connectivity for terrestrial users. In the context of NTN-aided communication, CD-RIS has shown great potential in improving the performance of cellular users with weak or no connectivity to the base station. For instance, as demonstrated in [14], CD-RIS aided UAVs (a specific use case of NTN) have clearly illustrated the usefulness of CD-RIS in NTN systems. Although CD-RIS aided NTN has shown significant potential in enhancing the performance of terrestrial users, the limitations associated with CD-RIS prevent us from fully benefiting from NTN. As discussed in the preceding sections, BD-RIS clearly surpasses CD-RIS by offering greater flexibility in reflecting signals in multiple controlled directions. This capability provides more extended coverage compared to CD-RIS and plays a significant role in increasing the performance of NTN.
V Case Study: BD-RIS aided NOMA LEO Satellite Communications
In this section, we provide a case study on BD-RIS aided NTNs. We will first discuss the considered system model and proposed optimization. Then, we provide the numerical results and their discussion.
Figure 4: (a) System model of NOMA BD-RIS aided LEO communications, (b) Achievable spectral efficiency of the system versus varying transmit power of LEO, (c) Achievable spectral efficiency of the system versus varying phase response elements.
@@ V-A System Model and Proposed Optimization
Let us consider a downlink scenario of BD-RIS aided satellite communication, as shown in Fig. 4 (a), where an LEO communicates with ground users (GUs) using the power domain NOMA. In practice, the GUs operate with low directional antennas, so closing links between LEO and GUs is challenging when the network operates on a high frequency. Moreover, the direct links between LEO and GUs might not be possible, considering the mobile scenario. To compensate for path losses due to poor link budget and improve the received signal strength, we consider fully-connected reflective BD-RIS installed on the optimal location and assisting the signal delivery from LEO to GUs. Thus, GUs can receive signals from direct and BD-RIS aided links. Assuming perfect acquisition of channel state information (CSI) by channel estimation techniques, this scenario provides an upper-bound solution. In addition, the GUs employ successive interference cancellation (SIC) to decode their information from the superimposed signal. In particular, the GU with a stronger channel condition can apply SIC to remove the other GU’s information before decoding its desired signal. Then, the GU with poor channel conditions cannot apply SIC and decode the signal directly by considering the other GU signal as noise.
This case study seeks to improve the achievable sum-rate of BD-RIS aided NOMA LEO satellite communication by optimizing the transmission power of LEO and PR of BD-RIS while ensuring the minimum data rate requirements of GUs. This is achieved by formulating a mathematical optimization problem for achievable sum-rate maximization which is subjected to several practical constraints such as minimum rate constraint, LEO power budget constraint, NOMA power allocation constraint, and BD-RIS PR constraint, respectively. The formulated optimization is nonlinear and NP-hard in nature due to coupling decision variables and NOMA interference. In practice, solving such problems is challenging due to high computational complexity. To make the optimization tractable, we first adopt the block descent coordinate method to split the joint optimization into two problems: 1) transmission power at LEO for a given PR at BD-RIS and 2) PR at BD-RIS, given the power allocation for LEO. Then, each problem is further transformed into linear form and simultaneously solved using a standard convex optimization such as CVX toolbox.